OpenStudy (anonymous):

What is the difference of 5x-2/4x - x-2/4x

OpenStudy (anonymous):

@passive13

OpenStudy (campbell_st):

well both fractions have the same denominator so can be written as \[\frac{5x - 2 -(x -2)}{4x}\] simplify the numerator

OpenStudy (anonymous):

How do I do that

OpenStudy (campbell_st):

well it would look like \[\frac{5x - 2 - x + 2}{4x}\] collect the like terms in the numerator

OpenStudy (campbell_st):

what do you get..when you collect the like terms...?

OpenStudy (anonymous):

6x-4/4x ?

OpenStudy (campbell_st):

ok... not quite right I'll write it this way \[\frac{(5 -1)x -2 + 2}{4x}\] try again

OpenStudy (campbell_st):

and that posting isn't quite right either

OpenStudy (campbell_st):

so what is 5 -1 equal to...?

OpenStudy (anonymous):

4x-4/4

OpenStudy (jhannybean):

yeah....that parenthesis. -_- \[\frac{ (5x-x)-2+2 }{ 4x }\]

OpenStudy (campbell_st):

getting closer 4x is correct so look at the 2nd part \[\frac{4x -2 + 2}{4x}\] what is -2 + 2 equal to

OpenStudy (anonymous):

0

OpenStudy (anonymous):

so the answer is 1?

OpenStudy (campbell_st):

ok... so now you have \[\frac{4x}{4x} \] thats correct...